How do I figure out The perimeter of a rectangular dog pen is 24 ft. The length is twice the width . What are the dimensions of the dog pen?
Answer by checkley71(8403)

PERIMETER=2*L+2*W OR 24=2(2W)+2W OR 24=4W+2W OR 24=6W OR W=4 LENGTH THEN L=2*4=8
PROOF 2*4+2*8=24 OR 8+16=24 OR 24=24